### v1

# Characterization of non-Markovianity with maximal extractable qubit-reservoir entanglement

Abstract
Understanding the dynamical behavior of a qubit in a reservoir is critical to applications in quantum technological protocols, ranging from quantum computation to quantum metrology. The effect of the reservoir depends on reservoir’s spectral structure, as well as on the qubit-reservoir coupling strength. We here propose a measure for quantifying the non-Markovian effect of a reservoir with a Lorentzian spectrum, based on the maximum qubit-reservoir quantum entanglement that can be extracted. Numerical simulation shows this entanglement exhibits a monotonous behavior in response to the variation of the coupling strength. We confirm the validity of this measure with an experiment, where a superconducting qubit is controllably coupled to a lossy resonator, which acts as a reservoir for the qubit. The experimental results illustrate the maximal extractable entanglement is progressively increased with the strengthening of the non-Markovianity.
pacs:
PACS number:
One of the most fundamental postulations of quantum mechanics is that the evolution of an isolated quantum system is governed by the Schrödinger equation. However, any system is inevitably coupled to its environment, which can be regarded as a reservoir containing many bosonic field modes 1 ; 2 ; 3 ; 4 ; 5 . According to the spectral structure, reservoirs can be classified into Markovian and non-Markovian. The Markovian reservoir has a flat spectrum, which makes it lose the memory about the past history. Consequently, the leakage of information and energy from the system to the environment becomes irreversible. Due to this irreversibility, the system’s dynamics, obtained by tracing over the reservoir modes, is governed by the Lindblad master equation 1 . The resulting time evolution displays a monotonous decaying feature, without any flow of information from the reservoir back to the system.
When the spectral structure of the reservoir is modified, it can become non-Markovian. Decaying resonators represent the most typical type of non-Markovian reservoirs, which have a Lorentzian spectrum 1 ; 2 . The memory effect of the reservoir depends on the ratio of the system-reservoir coupling to the spectral linewidth. When this ratio tends to infinity, the spectral spread can be neglected, so that the reservoir reduces to a single-mode resonator, which exchanges energy with the system in a reversible manner. In the opposite limit, the spectral spread can be considered infinite, so that the reservoir is approximately Markovian. In the intermediate regime, the reservoir displays some memory effect, which is gradually lost with the decrease of the coupling strength.
In the past decade, many efforts have been devoted to the quantification of non-Markovianity 6 ; 7 ; 8 ; 9 ; 10 . It has been shown that the quantum memory effect can be quantified by tracking the evolutions of two initially distinct quantum states in the reservoir 6 . When the reservoir is Markovian, the distinguishability between these two states is monotonously reduced and finally tends to zero when the information is completely leaked into the environment. For a non-Markovian process, this distinguishability displays an oscillatory pattern, as a consequence of the reversed information flow from the reservoir back to the system. Alternatively, the non-Markovianity can also be measured by using an ancilla that is initially entangled with the system in the reservoir, and tracking the ancilla-system entanglement during the system-reservoir interaction 7 . The reservoir’s memory effect is characterized by the extent to which the lost entanglement can be revived. Based on these measures, the non-Markovian effects have been experimentally explored in different systems, including photonic qubits 11 ; 12 ; 13 , nitrogen-vacancy centers 14 ; 15 ; 16 , and ion traps 17 . Recently, Gaikwad et al. explored the crossover between the Markovian and non-Markovian dynamics with a superconducting processor 18 , where the entanglement between a test qubit and an ancilla is used as a measure to quantify the non-Markovianity of the environment of the test qubit, comprised of a third qubit and its readout resonator.
We here propose an alternative measure for quantifying the memory effect for a general non-Markovian reservoir that interacts with a qubit, and present an experimental demonstration. Instead of introducing an ancilla that is initially entangled with the test qubit, the non-Markovianity is directly quantified by the extractable entanglement between the qubit and the reservoir generated by their interaction. When the qubit is initially in the excited state, the excitation is shared by the system and the reservoir due to their interaction. This excitation sharing results in the qubit-reservoir entanglement. The amount of the extractable maximum entanglement depends the non-Markovianity of the reservoir. In the strong coupling regime, the system-reservoir dynamics approximately reduces to the vacuum Rabi oscillation of the Jaynes-Cummings (JC) model, where the qubit-photon entanglement is completely extractable, with the maximum value approximating 1. With the decrease of the coupling strength, the amount of the extractable maximum entanglement is gradually reduced due to the spreading of the spectrum. In the Markovian limit, no entanglement can be extracted. The experimental data used to demonstrate the crossover between the Markovian and non-Markovian dynamics is taken from a previous experiment 19 , where a non-Hermitian JC model was synthesized with a superconducting qubit controllably coupled to its readout resonator. Here the readout resonator is taken as a reservoir. The experimental results confirm the theoretical predictions.
The system under consideration is composed of a qubit of frequency $\omega_{0}$ coupled to a non-Markovian reservoir whose spectrum is a continuum. The dynamics of the entire system, in the framework rotating at the qubit frequency, is governed by the Hamiltonian
$H=\left|g\right\rangle\left\langle e\right|\displaystyle\int\limits{}_{0}^{\infty}d\omega J(\omega)\lambda(\omega)e^{i\delta(\omega)t}a^{\dagger}(\omega)+H.c.,$ (1)
where $\left|e\right\rangle$ and $\left|g\right\rangle$ respectively denote the excited and ground state of the qubit, $J(\omega)$ is the spectral density of the reservoir, and $a^{\dagger}(\omega)$ and $a(\omega)$ are respectively the creation and annihilation operators for the bosonic mode with the frequency $\omega$ in the reservoir, which is coupled to the qubit with a coupling strength $\lambda(\omega)$ and a detuning $\delta(\omega)$. For a leaky resonator, the spectral shape is Lorentzian,
$J(\omega)=\frac{1}{\pi}\frac{\kappa/2}{(\omega-\omega_{0})^{2}+(\kappa/2)^{2}},$ (2)
where $\kappa$ represents the spectral line-width, which depends on the quality and the central frequency $\omega_{0}$ of the resonator. In conventional cavity quantum electrodynamics (QED) and circuit QED, cavities or resonators used to hold photonic modes can be modeled as such reservoirs 1 .
Such a non-Markovian reservoir is equivalent to the combination of a single lossy photonic mode of frequency $\omega_{0}$ and a Markovian reservoir, into which the photonic mode continuously leaks its energy with a rate $\kappa$. Hereafter, we will use the symbols “$Q$”, “$M_{0}$”, and “$R_{MR}$” to denote, respectively, the qubit, the photonic mode, and the Markovian reservoir, which corresponds to a flat continuum. When $Q$ is initially in the excited state $\left|e\right\rangle$ and $M_{0}$ in vacuum state $\left|0\right\rangle$, after an interaction time $t$, the qubit, the photonic mode, and the Markovian reservoir are evolved to the state
$\left|\Phi(t)\right\rangle=\left|\psi(t)\right\rangle\left|0\right\rangle_{MR}+\sqrt{1-\left\langle\psi(t)|\psi(t)\right\rangle}\left|g,0\right\rangle\left|1\right\rangle_{MR},$ (3)
where $\left|0\right\rangle_{MR}$ and $\left|1\right\rangle_{MR}$ denote the zero- and one-photon state for the Markovian reservoir, respectively. When $\xi=4\lambda_{0}/\kappa>1$, $\left|\psi(t)\right\rangle$ is given by
$\begin{split}\left|\psi(t)\right\rangle=\;&e^{-\kappa t/4}\{[\cos(\Omega t)+\frac{\kappa}{4\Omega}\sin(\Omega t)]\left|e,0\right\rangle\\ &-i\frac{\lambda_{0}}{\Omega}\sin(\Omega t)]\left|g,1\right\rangle\},\end{split}$ (4)
where $\Omega=\frac{\kappa}{4}\sqrt{\xi^{2}-1}$ with $\lambda_{0}=\lambda(\omega_{0})$. For $\xi<1$, the trigonometric functions are replaced by the corresponding hyperbolic Functions, with $\Omega=$ $\sqrt{\kappa^{2}/16-\lambda_{0}^{2}}$.
When the photon is leaked into the Markovian reservoir, it cannot be re-absorbed by the qubit and the information is completely lost. Once this event occurs, both $Q$ and $M_{0}$ jump to their ground state. The correlation between the Markovian reservoir and the qubit or the bosonic mode cannot be extracted. The part of quantum correlation that can be extracted is that associated with the non-unitarily evolved wavefunction $\left|\psi(t)\right\rangle$. This entanglement is characterized by the concurrence 20
${\cal C}=\left\langle\psi(t)|\psi(t)\right\rangle\sin(2\theta),$ (5)
where
$\theta=\arctan\left|\frac{4\lambda_{0}\sin(\Omega t)}{4\Omega\cos(\Omega t)+\kappa\sin(\Omega t)}\right|.$ (6)
The Concurrence ${\cal C}$ measures the unconditional entanglement between $Q$ and $M_{0}$, averaged over the trajectories with and without quantum jumps.
Figure 1: (color online). Theoretical results. a Extractable concurrence (${\cal C}$) between a qubit and a reservoir of photonic modes versus $\xi$ and $\tau$. The reservoir, initially in the vaccum state, is composed of photonic modes with a Lorentzian spectrum with a line width $\kappa$. The qubit, initially in the excited state, is on resonance with the central frequency of the continuum, $\omega_{0}$. Here $\tau=\kappa t/4$ denotes the rescaled interaction time, and $\xi$ is the ratio of the effective coupling strength between the qubit and the photonic mode with frequency $\omega_{0}$ to $\kappa/4$, and $\tau$. b Maximum of $\mathcal{C}$ during the evolution versus $\xi$. The dashed line is the derivative of $\mathcal{C}_{\max}$ with respect to $\xi$.
To illustrate the relation between the extractable quantum entanglement and the non-Markovianity of the system dynamics, we display ${\cal C}$ versus $\xi$ and the rescaled interaction time $\tau=\kappa t/4$ in Fig. 1a. When $\xi\rightarrow\infty$, the line width of the Lorentzian spectrum of the reservoir tends to 0 as compared to the qubit-reservoir coupling strength, so that the reservoir can be approximately regarded as a single photonic mode. Consequently, ${\cal C}$ nearly exhibits an oscillatory pattern with the period equal to half of that of the vacuum Rabi oscillation arising from the $Q$-$M_{0}$ swapping coupling, $\pi/\lambda_{0}$. ${\cal C}$ approaches the maximum value $1$ at the time $t=\pi/(4\lambda_{0})$, where the photon is equally shared by $Q$ and $M_{0}$. With the decrease of $\xi$, the line width of the reservoir cannot neglected. In this case, the photon swapping is only partly irreversible. Due to the progressive leakage of the information into $R_{MR}$, the local maximum of ${\cal C}$ gradually decreases. For $\xi\ll 1$, ${\cal C}$ almost remains to be 00, which can be interpreted as follows. In this weak coupling regime, the photon is immediately leaked to the Markovian reservoir as long as it is transferred from the qubit to the $M_{0}$ due to the strong dissipation. As a consequence, $M_{0}$ remains unpopulated, but mediates a coupling between $Q$ and $R_{MR}$, which results a completely irreversible dissipation process with a rate of $4\lambda_{0}^{2}/\kappa$.
To reveal how ${\cal C}$ can quantitatively manifest the extent of the non-Markovianity of the reservoir, we calculate the maximal value of ${\cal C}$ (${\cal C}_{\max}$) during the evolution for different values of $\xi$. For a given $\xi$, the time for the maximization of ${\cal C}$ is obtained by setting $d{\cal C}/dt=0$. The optimal time depends upon $\Omega(\xi)$ as
$t_{opt}=\left|\frac{2\arctan{\left(\frac{1}{2}\sqrt{\frac{\kappa^{2}+12\Omega^{2}-4\lambda_{0}\sqrt{\kappa^{2}+8\Omega^{2}}}{\Omega^{2}}}\right)}}{\Omega}\right|.$ (7)
Substituting $t_{opt}$ into Eq. (5), we obtain the maximally extractable concurrence, ${\cal C}_{\max}$. We show ${\cal C}_{\max}$ as a function of $\xi$ in Fig. 1b (solid line), which clearly shows that ${\cal C}_{\max}$ monotonously increases with $\xi$. The dashed line represents $d{\cal C}_{\max}/d\xi$ versus $\xi$, which illustrates that the changing rate of ${\cal C}_{\max}$ decreases with the increase of $\xi$, tending to 0 in the limit $\xi\rightarrow\infty$, where ${\cal C}_{\max}$ is saturated to 1. This implies that ${\cal C}_{\max}$ can serve as a measure for quantifying the non-Markovianity of the reservoir of bosonic modes with a continuum.
To confirm the validity of the approach, we use the data measured in a previous experiment focusing on the exceptional entanglement phase transition in a non-Hermitian qubit-photon model 19 . The experiment was performed with a circuit electrodynamics system, where an Xmon qubit with a negligible dissipative rate was coupled to its readout resonator with a non-negligible photonic decaying rate $\kappa=5$ MHz through a sideband interaction, mediated by a parametric modulation. Here the readout resonator is taken as the reservoir for the qubit, other than a part of the system. The effective qubit-photon coupling strength is controllable via the modulating amplitude. In the previous work we display the evolution of the qubit-photon entanglement associated with the no-jump trajectory, calculated with the output qubit-photon state projected onto the single-excitation subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle\right\}$. Such a conditional output state was obtained by reconstructing the qubit-photon density matrix with the assistance of an ancilla qubit, and then discarding the element associated to $\left|g,0\right\rangle$. The evolution of the post-selected output state coincides with the no-jump trajectory governed by the non-Hermitian Hamiltonian. On the contrary, as discussed above, the non-Markovianity is manifested here by the entanglement evolution associated with the whole output density matrix, reconstructed in the subspace $\left\{\left|e,0\right\rangle,\left|g,1\right\rangle,\left|g,0\right\rangle\right\}$.
Figure 2: (color online). Experimental results. a Measured ${\cal C}$ versus $\xi$ and $\tau$. The data is taken from a previous experiment 19 , which was performed in a circuit QED system, where the readout resonator of an Xmon qubit with a photonic decaying rate $\kappa$ is taken as the reservoir. b Measured ${\cal C}_{\max}$ versus $\xi$.
Fig. 2a presents the measured concurrence, without post-selection, as a function of $\xi$ and $t$. As expected, in the strong coupling limit $\xi\gg 1$, the unconditional qubit-photon concurrence displays nearly perfect periodic oscillations, which can be explained as follow. For a single photonic mode with a detuning $\delta(\omega)=\omega-\omega_{0}$, the oscillating frequency associated with the excitation swapping is $\chi_{\omega}=2\sqrt{\lambda^{2}(\omega)+\delta^{2}(\omega)/4}$. We here set $\left|\delta(\omega)\right|\ll\omega$ so that $\lambda(\omega)\simeq\lambda_{0}$. When the line-width of the Lorentzian spectrum of the resonator is much smaller than the coupling strength $\lambda_{0}$, i.e., $\left|\delta(\omega)\right|\ll\lambda_{0}$, we have $\chi_{\omega}\simeq 2\lambda_{0}$. This implies that all the bosonic modes within the spectral width of the reservoir are almost equally coupled to the qubit, keeping synchronous during interaction with the qubit, thereby collectively behaving like a single mode. With the decrease of $\xi$, the oscillatory feature is progressively weakened owing to the non-negligible spectral width.
Fig. 2b displays the maximum of the measured qubit-photon concurrence during the interaction, as a function of $\xi$. The slight fluctuations are mainly due to the imperfect control of the effective qubit-photon coupling strength. The results confirm that ${\cal C}_{\max}$ progressively increases with $\xi$, tending to 1 in the strong coupling regime. The monotonousness of ${\cal C}_{\max}$ implies that it can serve as a measure for characterizing the non-Markovianity.
In conclusion, we have shown that the maximal entanglement between a qubit and its reservoir extractable an be used as a measure for quantifying the non-Markovianity of the dynamics. When the qubit-reservoir coupling is much smaller than the spectral line-width of the reservoir, the leakage of the excitation from the qubit to the reservoir is almost irreversible. Consequently, most of the qubit-reservoir entanglement produced by their interaction is unextractable, reflecting the Markovian feature. With the increase of the coupling, the reservoir displays a non-Markovian effect, manifested by the nonvanishing qubit-reservoir entanglement that can be extracted. We illustrate this measure with a circuit QED experiment, where the readout resonator of a qubit acts as a reservoir of photonic modes with a continuous spectrum. By tuning the effective qubit-resonator coupling, we observed the progressive crossover from the Markovian to non-Markovian dynamics.
This work was supported by the National Natural Science Foundation of China (Grant No. 12274080, No. 11875108, No. 12174058, No. 12204105, No. 12374479), the NSF (Grant No. 2329027) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300200).